OpenStudy (anonymous):

Can someone explain to me how to find, how many solutions are in this equation, 2(x-3)=2x or other like it ???

Nnesha (nnesha):

distribute bracket by 2

Nnesha (nnesha):



OpenStudy (triciaal):

a*b = a*b on the left a= 2 (x-3) = b on the right a = 2 b = x b = b so (x- 3) = x x - 3 = x -3 is not = 0 no solution

OpenStudy (anonymous):